### v1

# Revisiting electron-capture decay for Galactic cosmic-ray data

Abstract
Electron-capture (EC) unstable species in Galactic cosmic rays constrain the time elapsed between nucleosynthesis and acceleration. They have also been advocated as tracers of reacceleration or gas inhomogeneities during their transport. The number of EC-unstable species grows with mass, with an expected EC-decay impact more important for larger atomic number and lower energy. We revisit the modelling of EC decay and its detectability in the context of recent unmodulated low-energy (Voyager) and high-precision data for heavy (AMS-02) and very-heavy nuclei (ACE-CRIS, CALET and Super-TIGER). We solve the transport equation for a multi-level configuration (up to any number of electrons attached) in the diffusion and leaky-box models. Their decayed fractions are found to be qualitatively similar but with very different absolute fluxes. We check that the standard two-level approximation, wherein the cosmic-ray nucleus is fully ionised or with one electron attached, is sufficient for most situations. We find that the impact of EC-decay is negligible in current data, except possibly for fluxes or ratios involving 51Cr, 55Fe, and Co. These conclusions are robust against significant uncertainties in the attachment and stripping cross-sections. This first analysis calls for further investigation, as several forthcoming projects (e.g., TIGERISS) are targeting $Z>30$ cosmic rays.
keywords:
Galactic cosmic rays , Propagation , Electron-capture decay , Electron attachment and stripping cross-sections
journal: Astroparticle Physics
[…]
1 Introduction
Most cosmic-ray (CR) data and studies focus on stable species, whose abundances and spectra are used to unveil their origin, acceleration, and transport (e.g., Strong et al., 2007; Tatischeff and Gabici, 2018). Unstable isotopes in CRs are rarer, but bring complementary information. The best known radioactive CRs are spontaneous $\beta$-unstable species with half-life at rest $t_{1/2}$ in the Myr range (and half-life in flight $\gamma t_{1/2}$, with $\gamma$ the Lorentz factor). These species are commonly referred to as CR clocks because, together with their associated ratios (e.g., 10Be and the 10Be/9Be ratio), they allow constraining the confinement time, or alternatively, the halo size of the Galaxy (Donato et al., 2002; Weinrich et al., 2020; Evoli et al., 2020; Maurin et al., 2022; Zhao et al., 2024). They are also sensitive to the presence of spatial inhomogeneities during their transport or in the surrounding gas distribution (Ptuskin et al., 1997; Ptuskin and Soutoul, 1998; Streitmatter and Stephens, 2001; Donato et al., 2002; Jacobs et al., 2023; De La Torre Luque and Linden, 2025).
Unlike spontaneous decay, electron-capture (EC) decay requires an electron from the innermost shell to interact within the nucleus to convert a proton into a neutron. EC decay in CRs is thus a multistep process with competing physical processes at play. Indeed, an electron must first be attached to the CR and then EC-decay must occur before the electron is stripped from the ion. Above GeV/n energies, electron stripping becomes more efficient than attachment: EC decay is effectively suppressed and CRs are fully ionised. This property can be used to estimate the time $t_{\rm elapsed}$ between nucleosynthesis and acceleration (Cassé, 1973; Cassé and Soutoul, 1975; Raisbeck et al., 1973). Freshly synthesised CRs at rest have available electrons and can undergo EC decay, whereas they are fully stripped and cannot decay via EC after the acceleration stage: consequently, the fraction of nuclei that have decayed via EC reflects the time they spent at low energy before acceleration. For instance, the EC decay of 57Co ($t_{1/2}=272$ days) measured in the associated elemental fluxes allowed to set $t_{\rm elapsed}\gtrsim{\cal O}(\rm yr)$ (Soutoul et al., 1978; Tueller et al., 1979; Dwyer and Meyer, 1985; Webber and Gupta, 1990), while the EC decay of 59Ni ($t_{1/2}=81$ kyr) in isotopic ratio data allowed to reach the lower limit $t_{\rm elapsed}\gtrsim 10^{4}-10^{5}$ yrs (Leske et al., 1992; Leske, 1993; Lukasiak et al., 1997; Wiedenbeck et al., 1999). This lower bound was recently nicely complemented by an upper bound $t_{\rm elapsed}\lesssim 2.6$ Myr, from the observations of a few $\beta$-unstable 60Fe ($t_{1/2}=2.62$ Myr) with the ACE-CRIS instrument (Binns et al., 2016).
EC-unstable species can also be used to trace energy changes during their journey, such as reacceleration in the Galaxy or energy losses during Solar modulation (Raisbeck et al., 1973). The idea here is to take advantage of the energy dependence of the attachment cross-section (the smaller the energy, the larger the cross-section). As a result, the probability of measuring a smaller (resp. larger) flux than expected, is linked to the energy gained (resp. lost) during the CR journey. For instance, the interpretation of measured ratios of daughter/parent EC-decay species (49Ti/49V and 51V/51Cr) initially concluded on the presence of a $100$ MeV/n reacceleration (Soutoul et al., 1998; Connell and Simpson, 1999; Connell, 2001a). This evidence was later disfavoured when accounting for the production cross-section uncertainties of these species (Jones et al., 2001a; Webber et al., 2003; Scott, 2005). These ratios are measured at the top-of-atmosphere (TOA), corresponding to interstellar (IS) energies $E_{\rm k/n}^{\rm IS}\approx E_{\rm k/n}^{\rm TOA}+\Delta E_{\rm k/n}$, with $\Delta E_{\rm k/n}$ an average loss in the Solar cavity (Caballero-Lopez and Moraal, 2004). With ACE-CRIS data taken at both solar minimum ($\Delta E_{\rm k/n}\sim 200$ MeV/n) and maximum ($\Delta E_{\rm k/n}\sim 350$ MeV/n), Niebur et al. (2003) argued for the detection of the energy-dependent EC decay. The same data were also used to set indirect constraints on the CR low-energy IS flux shape (Mewaldt et al., 2004; Caballero-Lopez et al., 2007), before direct measurements outside the Solar cavity became available (Stone et al., 2013, 2019; Cummings et al., 2016).
EC-decay species were also invoked to trace possible inhomogeneities in the IS medium (ISM) where CR travel, as ions cannot pick up electrons in depleted zones (Letaw et al., 1985), therefore suppressing EC decay. This could trace for instance the time of our last passage in the spiral arms (Benyamin et al., 2017), or the ionic state of nuclei could be sensitive to the very local inhomogeneities (Tylka et al., 1995a). There is also an interesting link with $\gamma$-ray spectroscopy (Diehl et al., 2006): pure EC-decay species, whose daughter is a $\gamma$-ray emitter, enable to inspect the deduced Galactic CR (GCR) source abundance and direct production of 44Ti (Benyamin and Shaviv, 2018). However, beside the last two studies, almost no work was published on EC-decay species in GCRs in the last 20 years. Interested readers are referred to the reviews of that time for more details (Letaw et al., 1985; Mewaldt et al., 2001).
The interest in EC-unstable CR species should be renewed soon, though. Indeed, a wealth of experiments are planned to explore, above a few hundreds of MeV/n, species with atomic number $Z>30$, i.e., in a region with numerous EC-unstable species. Actually, some of Super-TIGER data are still being analysed (Sasaki, 2021; Rauch et al., 2023), while the TIGER–ISS is scheduled for launch in 2027 (Rauch et al., 2024), and the NUCLEON-2 (Bulatov et al., 2019; Vasil’ev et al., 2021) and HERO (Karmanov et al., 2019; Kurganov et al., 2023; Podorozhny et al., 2024) satellite missions are envisioned in the next decade (see, for instance Fig. 8 of Maurin et al. 2025). Besides, ongoing experiments provided measurements with unprecedented precision, in regimes where the impact of EC decay could be directly detected, including, at GeV/n TOA energies, elemental fluxes measured up to Fe at the few percent-level precision by AMS-02 (Aguilar et al., 2021b), elemental ratios $26\leq Z\leq 40$ measured by Super-TIGER (Murphy et al., 2016) and $14\leq Z\leq 44$ by CALET (Adriani et al., 2025), isotopic ratios $29<Z<38$ by ACE-CRIS at a few hundred of MeV/n (Binns et al., 2022), and last but not least, IS fluxes up to Ni by Voyager down to a few tens of MeV/n (Cummings et al., 2016). To take a broader view, we recall that EC decay during propagation impacts the derived GCR source abundances (Letaw et al., 1985), and that the latter are key to connect to the various nucleosynthesis production processes leading to ultra-heavy nuclei (Arnould and Goriely, 2003; Arnould et al., 2007; Thielemann et al., 2017; Arcones and Thielemann, 2023).
On the modelling side, EC decay has not been discussed much either, recently, with the data interpreted in the leaky-box or weighted-slab approximation. To track EC-decay species, these models relied on a 2-level description of the transport equation, with a fully stripped and a hydrogen-like ion state only (Letaw et al., 1984a). The same 2-level description is also used in the diffusion model GALPROP for $Z\leq 30$ species (Porter et al., 2022). In contrast, EC decay is not included in the DRAGON code (Evoli et al., 2018), the authors arguing that EC decay has no impact on the calculated fluxes (which they checked by running the GALPROP code).
For all these reasons, it is important to revisit the modelling and impact of EC decay for GCR fluxes, especially in the context of recent $Z\leq 40$ and forthcoming $Z>30$ high-precision data. In Sect. 2, we recall the GCR transport equation and the specifics of EC decay, highlighting the various timescales involved. In Sect. 3, we present our results concerning the generic impact of EC decay and the validity of the two-level ion approximation. In Sect. 4, we discuss the detectability prospects of EC decay in current GCR isotopic and elemental data. In Sect. 5, we conclude and draw the next steps of this study. For the sake of readability, we postponed the most technical details in Appendices: A recaps the formulae for single electron attachment and stripping in the K-shell and their generalisation; B provide analytical solutions for the differential density of EC-unstable species considering any level of attached electrons, for both the leaky-box model (hereafter, LBM) and 1D diffusion model (hereafter, 1D-DM).
2 Transport equation and timescales
GCR transport is described by a diffusion-advection equation, which includes spatial transport, momentum transport, and various source and sink terms.
2.1 From differential density to modulated fluxes (without EC)
The differential density $N_{i}$ of a stable or $\beta$-unstable species $i$ as a function of energy $E$ is given by (e.g., Berezinskii et al., 1990)
$\begin{split}-\vec{\nabla}\left[D(E)\vec{\nabla}N_{j}-\vec{V}_{c}N_{i}\right]+\frac{\partial}{\partial E}\left[b_{\rm tot}(E)N_{i}-(v/c)^{2}K_{pp}\frac{\partial N_{i}}{\partial E}\right]\\ +\left(\Gamma^{\rm inel}_{i}+\Gamma^{\beta}_{i}\right)N_{i}=q_{i}+\sum_{j>i}\Gamma^{\rm prod}_{j\to i}N_{j}+\sum_{k>i}\Gamma^{\beta}_{k\to i}N_{k}\,.\end{split}$ (1)
In the first line, from left to right, the transport terms are: spatial diffusion $D$ (CR scattering off the magnetic turbulence), convective transport $V_{c}$, energy losses (ionisation, Coulomb, and adiabatic) and gains (momentum diffusion $K_{pp}$). In the second line, the CR destruction (left-hand side) and production from heavier species (right-hand side) involve nuclear cross-sections
$\Gamma^{\rm inel}_{i}=\sum_{t\,\in\,\rm ISM}\,n_{t}\,v_{i}\,\sigma_{i+t}^{\rm inel}\quad{\rm and}\quad\displaystyle\Gamma^{\rm prod}_{j\to i}=\sum_{t\,\in\,\rm ISM}\,n_{t}\,v_{i}\,\sigma_{j+t\to i}^{\rm prod}\,$, (2)
and $\beta$ decay (half-life $t_{1/2}$) as a sink (decay of $i$) or a source (decay of CR $k$ into $i$),
$\Gamma^{\beta}=\frac{\ln(2)}{\gamma t_{1/2}}\,.$ (3)
Finally, the injection rate $q_{i}$ describes the primary production.
This transport equation can be solved by different approaches and associated codes: numerically after discretisation on spatial and energy coordinates – e.g., DRAGON (Evoli et al., 2018), GALPROP (Porter et al., 2022) and PICARD (Kissmann, 2014) –, via stochastic differential equation based Monte Carlo simulations (e.g., CRPROPA, Merten and Aerdker 2025), or from semi-analytical solutions in simplified geometries based on Green functions, series expansion, etc. (e.g., the LBM and 1D-DM in USINE [^1] https://dmaurin.gitlab.io/USINE/, Maurin 2020). Once $N$ is calculated for all CR isotopes, the (nearly isotropic) CR interstellar (IS) flux is given by $\psi^{\rm IS}=N(v/4\pi)$. To compare the modelled fluxes to the measured ones, Solar modulated top-of-atmosphere (TOA) fluxes are calculated from IS ones, using the force-field approximation (Gleeson and Axford, 1967, 1968). The energy is transformed according to
$E_{\rm k/n}^{\rm TOA}=E_{\rm k/n}^{\rm IS}-\phi_{\rm FF}\times(Z/A)\,$, (4)
with $Z$ and $A$ the atomic and mass number of the species, $\phi_{\rm FF}$ the solar modulation potential varying from $\sim 500$ MV (Solar minimum) to 1200 MV (Solar maximum). The flux at the top of the atmosphere is related to the interstellar flux by $\psi^{\rm TOA}(E^{\rm TOA})=\psi^{\rm IS}(E^{\rm IS})\times(p^{\rm TOA}/p^{\rm IS})^{2}$. In the above equations, $p$ is the momentum, $E$ is the total energy, and $E_{\rm k/n}=E_{\rm k}/A$ is the kinetic energy per nucleon, with $E_{\rm k}=E-m$; the rigidity, $R=pc/Ze$, is also used later on in the paper.
2.2 Specifics of EC decay in GCRs
In spontaneous $\beta$ decay, a neutron (resp. proton) is converted into a proton (resp. neutron) in a nucleus $X$,
$^{A}_{Z}X\;\stackrel{{\scriptstyle\beta^{-}}}{{\to}}\;^{\;\;\;\,A}_{Z+1}X+e^{-}+\overline{\nu_{e}}\,$,
$^{A}_{Z}X\;\stackrel{{\scriptstyle\beta^{+}}}{{\to}}\;^{\;\;\;\,A}_{Z-1}X+e^{+}+\nu_{e}\,.$
In contrast, EC decay is a process in which the nucleus captures one of its atomic electrons,
$^{A}_{Z}X+e^{-}\;\stackrel{{\scriptstyle EC}}{{\to}}\;^{\;\;\;\,A}_{Z-1}X+\nu_{e}\,$,
usually from the innermost shell ($K$-shell), which has the highest probability to interact with the nucleus. As a result, EC decay is possible only if it is preceded by electron attachment, and only if electron stripping is not faster than the decay time. This additional decay channel thus results from a balance (or competition) between the three rates at play ($\Gamma^{\rm att}$, $\Gamma^{\rm str}$ and $\Gamma^{\rm EC}$), where the first two processes occur in the ISM (where electrons are available), while the last one can occur anywhere. For a CR species $i$, the first two rates are defined as
$\Gamma^{\rm att}_{i}=\sum_{t\,\in\,\rm ISM}\,n_{t}\,v_{i}\,\sigma^{\rm att}_{i+t}\quad{\rm and}\quad\Gamma^{\rm str}_{i}=\sum_{t\,\in\,\rm ISM}n_{t}\,v_{i}\,\sigma^{\rm str}_{i+t}\,$, (5)
where $\sigma^{\rm att}$ and $\sigma^{\rm str}$ are the cross-sections for the attachment and stripping of electrons in the ISM (see A).
As a consequence of the energy and mass dependence of the latter quantities, ions in GCRs are usually fully-ionised. Indeed, stripping is dominant over a few tens of MeV/n and for light nuclei (see Sect 2.3), so that EC decay is suppressed. This explains why EC decay can be used to measure the time elapsed between nucleosynthesis and acceleration: species at rest in the ISM have electrons attached (EC decay enabled), but lose them once they are accelerated (EC decay blocked). Furthermore, if reacceleration occurs during the transport, the bulk of GCRs can gain a few hundreds of MeV/n, moving from a situation where EC decay is favoured (electron attachment is possible), to one where it is disfavoured (ions are mostly fully ionised).
EC decay is the sole decay mode in proton-rich nuclei, if the binding energy difference between the initial and final states is below $2m_{e}\simeq 1022$ keV. Otherwise, EC and $\beta^{+}$ are competing processes, although their respective branching ratios are not always experimentally available (Kondev et al., 2021). Table LABEL:tab:EC_clocks lists all nuclei undergoing pure EC decay as extracted from Nubase (Kondev et al., 2021), along with their half-lives, final stable daughters and, where available, their multistep decay chains.
The modification of the transport Eq. (1) to account for EC decay, along with the explicit solutions for the LBM and 1D-DM are detailed in B.
2.3 Dominant processes and timescales
We show in this section the timescales of the processes involved in 1D-DM, consisting of a thin disc of half-width $h$ (source and gas) and a diffusive halo of half-width $L$ (in an infinite plan). The interplay between these timescales shapes the isotopic and elemental GCR fluxes. The escape time from the diffusive volume (or diffusion time) is $t_{\rm diff}=L^{2}/(2D)$. The inelastic, attachment and stripping times in the gaseous disc are obtained from the inverse of the rates defined in Eqs. (2) and (5); the decay time is the inverse of the decay rate Eq. (3).
Table 1: Typical energy and CR charge (or atomic number) dependence of timescales in the 1D-DM (see B) in the relativistic regime ($\beta\approx 1$): diffusion (diff), inelastic interactions (inel), electron radiative attachment (ra) and non-radiative attachment (nra), electron stripping (str), and $\beta$ decay; see A for a full discussion of the attachment and stripping cross-sections.
Rates
$t_{\rm diff}$ $\propto E^{-0.5}$
$t_{\rm inel}$ $\propto A^{-2/3}$
$t_{\rm att}$ $\propto E\,Z^{-5}$ (ra)
$\propto E\,Z^{-5}$ (nra)
$t_{\rm str}$ $\propto Z^{2}/\ln E$
$t_{\rm EC}$ $\propto E$
Figure 1: Timescales for the processes listed in Table 1, as a function of the atomic number $Z$, for three energies (from left to right). The attachment time (pink dotted line) is broken down into the radiative (thin-dotted grey line) and non-radiative (thick-dotted grey line) contributions, see A.1. The black dots represent the effective decay time ($\gamma\,\tau_{\rm EC}$) for each EC-unstable isotopes reported in Table LABEL:tab:EC_clocks (highlighted species are compared to CR data in Sect. 4).
Energy range, transport and destruction timescales
We gather in Table 1 the most salient energy or atomic number dependence of these timescales. In Fig. 1, we show them as a function of the atomic number $Z$, for three energies. We choose the minimum energy 10 MeV/n (left panel) as it corresponds to the lowest IS energies measured by CR experiments (see Sect. 4), and 1 GeV/n as the maximum energy (right panel), because EC decay is blocked above [^2] In passing, this is a striking difference between EC and spontaneous decay, as the former is blocked when $t_{\rm att}\gg\min(t_{\rm diff},\,t_{\rm inel})$, whereas the latter is blocked when $\gamma t_{1/2}\gg t_{\rm diff}$.. As already reported in the literature (e.g., Derome et al., 2019), below a few GeV/n, destruction (blue dashed line) is significant over diffusion (green solid line) for heavy nuclei (i.e., $t_{\rm inel}<t_{\rm diff}$ for $Z\gtrsim 10$). Actually, as discussed in Vecchi et al. (2022), diffusion prevails at increasingly larger energies for growing CR masses.
Attachment and stripping timescales
The total electron attachment time (pink dotted line) is dominated by radiative attachment (thin grey dotted line): the non-radiative one (thick grey dotted line) is negligible, except for light nuclei below tens of MeV/n (left panel). Attachment is more efficient for heavy species ($t_{\rm att}\propto Z^{-5}$), while this is the opposite for stripping ($t_{\rm str}\propto Z^{2}$), see Table 1. Attachment is also more efficient at low (10 MeV/n) than at high energy (1 GeV/n), whereas the energy dependence for stripping is milder. Actually, in the sub-GeV range, the energy dependence is non-trivial for these two processes. Overall, we have $t_{\rm att}<t_{\rm str}$ for heavy nuclei and low energy only, which is the optimal regime to ensure at least one electron can be attached.
Decay time
EC decay is possible if an electron is attached, that is, if $t_{\rm att}<\min(t_{\rm diff},\,t_{\rm inel})$, which is the case at low energy and for heavy nuclei. The majority of the known EC isotopes have $t_{\rm EC}\ll t_{\rm str}$ (black dots below the orange dash-dotted line). For these short-lived CRs, decay occurs immediately after attachment. For the remaining CRs (i.e., with $t_{\rm EC}\gtrsim 10^{-4}$ Myr), the interplay among stripping, attachment and decay times has to be taken fully into account to determine their effective decay rate. We refer readers to Letaw et al. (1984a, b, 1985) for a complementary analysis of these behaviours (carried out in the LBM).
3 Impact of EC-decay, uncertainties, and validity
Throughout this section, our reference calculation (with EC decay) is a 2-level model, denoted $N_{2\penalty 10000\ {\rm lev}}$, as considered in the literature (i.e., either fully ionised CR or with a single electron attached). We use the 1D-DM and the best-fit transport parameters of Ferronato Bueno et al. (2024), where neither convection nor reacceleration is favoured by CR data.
3.1 Preamble
For this analysis, we neglect energy losses, which allows us to use the simple analytical solutions reported in B. Energy losses shape GCR spectra below hundreds of MeV/n, roughly shifting downwards their energy: GCRs at $E_{\rm k/n}$ actually come from $E^{\prime}_{\rm k/n}>E_{\rm k/n}$, that is, an energy where EC decay is less efficient and impacting (see previous section). Consequently, the no-energy loss calculation overestimates the impact of EC decay. We explore below several aspects of EC decay: validity of some approximations, impact of uncertainties, detectability of EC decay in current CR data (Sect. 4). We will show that our conclusions, obtained in the no-energy loss framework, are conservative, that is, they would be even more valid if energy losses were to be accounted for.
3.2 Impact of EC decay on isotopic IS fluxes
Figure 2: Colour-coded EC-decayed fraction, Eq. (6), in the $t_{1/2}$$Z$ plan for the 1D-DM (top) and LBM (middle), for three IS energies. The two near-vertical lines separate short-lived, intermediate-lived, and long-lived isotopes (see text for details). Symbols indicate existing EC-unstable isotopes (see Table LABEL:tab:EC_clocks), highlighting the names of those for which CR data exist (see Sect. 4).
Impact in the 1D-DM
The top panels of Fig. 2 show the decayed fraction,
$f=1-\frac{N_{2\penalty 10000\ {\rm lev}}}{N_{1\penalty 10000\ {\rm lev}}}=1-\frac{\psi_{2\penalty 10000\ {\rm lev}}}{\psi_{1{\rm lev}}}\,$, (6)
as a function of a generic $t_{1/2}$ and $Z$. In the above equation, $N_{2\penalty 10000\ {\rm lev}}$, defined in (19), corresponds to the calculation tracking only two states (fully ionised, and single electron attached), as used in the literature, whereas $N_{1\penalty 10000\ {\rm lev}}$ is the calculation without EC decay (only the fully ionised state is enabled).
As expected from the timescales, EC decay has a maximum impact on isotopic fluxes at low energy (left panel), and mostly at high $Z$ and short decay times (top-left corners in the various panels). We stress that above a few GeV/n (not shown), the decayed fraction becomes less than 1% for EC-unstable species currently measured in CR data (those with their element name in the plot); above tens of GeV/n, this applies to all EC-unstable isotopes.
The near-vertical line on the right-hand side of all panels corresponds to $\gamma\,\tau_{\rm EC}=10\,t_{\rm diff}$: isotopes on the right-hand side of this line mostly escape from the Galaxy before having time to decay, so that their decayed fraction $f<1\%$. The other near-vertical line (on the left) corresponds to the condition $\gamma\tau_{\rm EC}=0.1\,t_{\rm str}$: isotopes on the left-hand side of this line (i.e., where most EC-decay isotopes lie) decay immediately after they attach an electron, hence, their decay is solely controlled by the attachment time. As a result, the decayed fraction is $f\simeq 100\%$ for all species at low energy (left panel), whereas at higher energy (central and right panel), it goes from a few percent up to 100% for growing $Z$, following the behaviour of the attachment time. In the in-between region (i.e., intermediate-lived isotopes), the decay impact depends on the interplay between the EC decay, stripping, and attachment times, and can go from 100% to less than 1%, depending on the energy and the charge of the isotope.
Figure 3: Same axes and configuration as in Fig. 2, but now the colour scale shows the ratio of the LBM to the 1D-DM calculation, defined in Eq. (8). The size of the circles encodes the decayed fraction $f$ of the EC-unstable isotopes.
Differences between the LBM and 1D-DM
Studies on $Z>30$ species have mostly been carried out in the context of LBM, so we extend our discussion to this model. We first stress that, for stable species, the solution of the LBM and the 1D-DM are formally similar, provided that (e.g., Jones et al., 2001b)
$\tau_{\rm esc}=\frac{hL}{D}\,.$ (7)
We use the above relation in our calculations, in order to have meaningful comparisons between the two models. However, the 1D-DM is not reducible to the homogeneous LBM for $\beta$-unstable species (Prishchep and Ptuskin, 1975), and we expect the same irreducibility for EC-unstable species. Indeed, this non-equivalence is directly seen from the different formal solutions, Eqs. (31) for the LBM and (32) for the 1D-DM, where the LBM solution is more dominated by the decay term than is the 1D-DM one.
The decayed fraction in the LBM is shown in the bottom panel of Fig. 2. Qualitatively, the trend is very similar in the 1D-DM and LBM (compare the top and middle panels). However, quantitatively, the decayed fraction shows significant differences for intermediate-lived isotopes (those between the two vertical lines). A complementary view is provided by forming the ratio of the LBM to the 1D-DM solutions (Fig. 3),
${\cal R}^{\rm LBM/1D-DM}=\frac{N_{2\,{\rm lev}}^{\rm LBM}}{N_{2\,{\rm lev}}^{\rm 1D-DM}}=\frac{\psi_{2\,{\rm lev}}^{\rm LBM}}{\psi_{2\,{\rm lev}}^{\rm 1D-DM}}\,.$ (8)
For long-lived isotopes (close or beyond the vertical line on the right), the ratio is $1$, as expected for stable species. For intermediate-lived species with $Z>40$ and $E_{\rm k/n}<100$ MeV/n, the LBM is a factor 1.5 larger than the 1D-DM calculation, whereas just at the transition between short-lived and intermediate-lived species (near-vertical line on the left-hand side of the plot), the LBM can be half the value of the 1D-DM, or even less for $Z<30$ at very low energy (bottom left corner of the left panel). Among the currently measured CR fluxes, $^{44}{\rm Ti}$ is the EC-stable species most impacted by the model choice.
Figure 4: Same as bottom panel of Fig. 2, but for the ratio ${\cal R}^{\sigma_{\rm new}}$, see Eq. (9), when halving the attachment (top) and doubling the stripping cross-sections (bottom).
3.3 Impact of $\sigma^{\rm att}$ and $\sigma^{\rm str}$ uncertainties
As discussed in A, the attachment and stripping cross-sections ($\sigma^{\rm att}$ and $\sigma^{\rm str}$) used in the GCR literature are based on data and parametrisations dating back from the 1970s (Wilson, 1978; Crawford, 1979; Letaw et al., 1984b). Moreover, these formulas are derived in the context of single electron attachment, and their extrapolation to higher ionised states could introduce additional uncertainties in the models (see A.3). Therefore, it is important to quantify the impact of a variation in $\sigma^{\rm att}$ and $\sigma^{\rm str}$ on the isotopic fluxes, and for that, we focus on the ratio (in the 2-level configuration)
${\cal R}^{\sigma_{\rm new}}=\frac{N_{2\,{\rm lev}}^{\sigma_{\rm new}}}{N_{2\,{\rm lev}}^{\sigma_{\rm ref}}}=\frac{\psi_{2\,{\rm lev}}^{\sigma_{\rm new}}}{\psi_{2\,{\rm lev}}^{\sigma_{\rm ref}}}\,.$ (9)
In Fig. 4, this ratio is shown for a halved $\sigma^{\rm att}$ (top) and a doubled $\sigma^{\rm str}$ (bottom). This choice allows us to show the two cases with the same colour scale, easing a visual comparison and discussion (of the impact of a 50% uncertainty on these cross-sections). Indeed, both changes lead to less decay, hence a larger $N^{\sigma_{\rm new}}$ and ${\cal R}^{\sigma_{\rm new}}>1$, as it takes a longer (resp. shorter) time to attach (resp. strip) an electron, slowing down and decreasing the decay probability. We stress that if, instead, we were to double $\sigma^{\rm att}$ and half $\sigma^{\rm str}$, this would lead to the opposite behaviour – more decay, smaller $N^{\sigma_{\rm new}}$, and ${\cal R}^{\sigma_{\rm new}}<1$ –, while following a similar pattern in terms of the dependence with $E_{\rm k/n}$, $t_{1/2}$ and $Z$.
Uncertainties on $\sigma^{\rm att}$
In the top panel, we observe a growing impact of EC decay towards lower energies (i.e., from right to left panels), and going from intermediate- to short-lived isotopes for larger $Z$ (upper-left corners of the plots). Long-lived isotopes are insensitive to the modified cross-section, i.e., ${\cal R}^{0.5\sigma^{\rm att}}\!(t_{1/2}\gtrsim 1\,{\rm Myr})\!=\!1$. The impact of halving $\sigma^{\rm att}$ (top panel) is important for short-lived nuclei (${\cal R}^{0.5\sigma^{\rm att}}\!\!=2$), as attachment directly drives the decay time. The region of intermediate-lived nuclei – near the vertical line on the left-hand side – is also impacted, as it is very sensitive to the balance between attachment, stripping, and decay. Looking back at the timescales in Fig. 1, among the species with CR data (i.e., highlighted names in the figure), the curve for $t_{\rm att}$ with a halved decay time would get further away from 53Mn and 41Ca, but closer to 44Ti, 55Fe, etc. This reflects into a maximum impact for the latter, where the decayed flux is changed by more than a factor two at low energy (top left panel in Fig. 4).
Uncertainties on $\sigma^{\rm str}$
A comparison of the bottom and top panels of Fig. 4 shows that changing the stripping and attachment times have very different impacts on the decaying fractions. In addition to have no effect on long-lived isotopes and high energies, there is also no effect on short-lived nuclei, as their decay is not limited by stripping (left side of the plots). Intermediate-lived nuclei are impacted at the same level and for the same nuclei as when halving $\sigma^{\rm att}$ (compare the top and bottom panels), owing to a similar modification of the balance between attachment, stripping, and decay. This impact is ${\cal R}^{2\sigma^{\rm str}}\!\!=2$ at 10 MeV/n (left panel). It decreases at higher energies (right panels), while the most impacted nuclei are shifted towards larger $Z$ and $t_{1/2}$.
Overall impact of these uncertainties
A last step must be taken to conclude whether more precise cross-sections are needed or not for $\sigma^{\rm att}$ ad $\sigma^{\rm str}$. Indeed, the impact of cross-section uncertainties (${\cal R}^{\sigma_{\rm new}}$) must be considered in the light of the overall EC-decayed fraction. For existing isotopes, the latter is indicated in Fig. 4 by the size of the circles (or, equivalently, encoded in the $f$ values shown in the top panels of Fig. 2):
  • 1.
    in the bottom-right region of the $t_{1/2}-Z$ plan, we have ${\cal R}^{\sigma^{\rm att}_{\rm new}}\simeq{\cal R}^{\sigma^{\rm str}_{\rm new}}\simeq 1$ and $f\lesssim 1\%$. As there is no foreseeable scenario in which CR data will reach the percent precision on isotopes, large uncertainties on $\sigma^{\rm att}$ and $\sigma^{\rm str}$ (up to a factor two or more) are non-impacting.
  • 2.
    in the top-left region of the $t_{1/2}-Z$ plan, we have ${\cal R}^{\sigma^{\rm att}_{\rm new}}\gtrsim 2$, ${\cal R}^{\sigma^{\rm str}_{\rm new}}\approx 1$ and $f\sim 100\%$. The latter number means that EC-decay decreases the fluxes by a factor $\gtrsim 100$ and is insensitive to uncertainties on $\sigma^{\rm str}$, but that a 50% uncertainty on $\sigma^{\rm att}$ translates to a factor of two uncertainty on this suppression factor: while significant, this does not qualitatively change the conclusions regarding the detectability of EC-decay in CR data (the data precision goes from a few percent to tens of percent, see Sect. 4).
  • 3.
    intermediate-lives nuclei (left-hand side of the near-vertical line) with $Z\approx 30$ (i.e., 44Ti, 55Fe, etc.): this is a region, at low-energy, where both ${\cal R}^{\sigma^{\rm att}_{\rm new}}$ and ${\cal R}^{\sigma^{\rm str}_{\rm new}}$ are $\gtrsim 2$, and $5\lesssim f\lesssim 50\%$ (depending on the energy). Consequently, the cross-section uncertainties maximally impact the conclusions for them: the EC-decay impact can go from very small (few percent) to very significant (up to a factor 2), compared to the CR data precision.
We stress that we obtain similar results for the LBM, and also for multi-level models (not shown). We will come back to these uncertainties in Sect. 4. However, we can already say that, overall, most EC-unstable isotopes are insensitive to large uncertainties on $\sigma^{\rm str}$, while their behaviour is only qualitatively insensitive to uncertainties on $\sigma^{\rm att}$. Only the flux of isotopes $t_{1/2}\approx{\cal O}(\rm kyr)$ at IS energies below a few hundreds of MeV/n is critically sensitive to uncertainties on both cross-sections.
Figure 5: Same as in previous figures, but showing $n$ at which convergence is reached (top row), and the relative difference between the literature approximation $n=2$ and the converged solution (bottom row); both calculations are for the 1D-DM.
3.4 Validity of currently used approximations
Previous studies (Letaw et al., 1985; Soutoul et al., 1998; Connell, 2001a; Jones et al., 2001a; Webber et al., 2003; Scott, 2005) rely on two assumptions. First, partially ionised nuclei have different charges, so that all CR species (stable ones included) should be differently affected by diffusion. These works do not explicitly state whether they account for this. Second, GCR calculations do not consider more than one electron attached, because the corresponding populations drop quickly (Crawford, 1979; Letaw et al., 1984b). Experimentally, various ionic states, up to hundreds of MeV/n, have been observed in the Solar neighbourhood. A GCR origin for high-ionic Fe states was even briefly considered (Biswas et al., 1990; Tylka et al., 1995a). However, the presence of ionic states is only firmly established for transient Solar energetic particles (Tylka et al., 1995b; Klecker et al., 2007; Reames, 2018) and anomalous CRs (Dutta et al., 1993; Mewaldt et al., 1996; Baranov et al., 2007).
Accounting for an effective charge in the transport
Different ionised states of the same isotope have different charges. As diffusion depends on the rigidity, one should in principle account for this charge difference. We checked that the solution with and without this approximation differs by $\lesssim 1\%$ at all energies and for all species (not shown). This validates the use of the same charge for the transport of all ionic states of the same isotope.
Multi- vs 2-level calculation
As defined in Eq. (19) and B.1, the calculation with $n$ levels considers CR ions with zero, one, two, etc., up to n1n-1 electrons attached. We define the quantity $\Delta_{(n)}$,
$\Delta_{(n)}=\left|\frac{N_{n\,{\rm lev}}}{N_{(n-1)\,{\rm lev}}}-1\right|\,$,
to track the relative difference between the calculation carried out up with $n$ and n1n-1 levels. We consider to have reached the asymptotic solution (or convergence) at $n$ if $\Delta_{(n)}\leq 1\%$.
The top panel of Fig. 5 shows $n$ at which the convergence is reached: above a few hundreds of MeV/n (IS), the 2-level approximation is sufficient for all EC-unstable species, but for a few nuclei with $Z\gtrsim 40$ and $t_{1/2}\gtrsim 10$ kyr. Nevertheless, at 10 MeV/n, the convergence is reached for $n=13$ levels for intermediate-lived nuclei of charge Z=3040Z=30-40. For the LBM, similar conclusions are reached (same number of levels to reach convergence), though with slightly tighter contours along the $t_{1/2}$ axis.
The bottom panel shows the relative difference between the converged solution (assimilated to be $N_{13\,{\rm lev}}$) and $N_{2\,\rm lev}$. The flux is further suppressed when extra levels are present, and the largest for regions of large $n$ (compare the top and bottom panels). The relative difference between $N_{3\,\rm lev}$ and $N_{2\,\rm lev}$ (not shown) indicates that a large fraction of the extra-suppression is already contained in the $n=3$ level calculation, whereas each new level has a decreasing impact. This is not surprising, because going from 1 to 2 attached electrons doubles the EC-decay rate (two instead of one electron are available in the $K$-shell), whereas supplementary electrons (in other shells) do not enhance the decay and come at the cost of multiplying the overall attachment time.
Among the isotopes for which CR data exist (i.e., those with the names highlighted in the plot), the converged solution differs at most by $\lesssim 20\%$, and only for very low energy (bottom left panel). More importantly, in the range targeted by TOA current experiments, that is, above several hundreds of MeV/n in equivalent IS energies (in-between the middle and right panels), the difference in the decayed fraction does not even reach the percent level, validating the use of the 2-level approximation. Similar conclusions are reached for the LBM.
Table 2: CR isotopic fraction (weighted mean and weighted error) extracted from the following TOA data in CRDB: $^{7}{\rm Be}$/Be (Hagen et al., 1977; Garcia-Munoz et al., 1977; Webber et al., 1977; Buffington et al., 1978; Webber and Kish, 1979; Wiedenbeck and Greiner, 1980; Garcia-Munoz et al., 1981; Connell, 1998, 2001b; Nozzoli and Cernetti, 2021), $^{37}{\rm Ar}$/Ar (Young et al., 1981; Webber et al., 1985; Ogliore et al., 2009), $^{41}{\rm Ca}$/Ca (Young et al., 1981; Webber et al., 1985; Leske and Wiedenbeck, 1993; Connell, 2001b; Ogliore et al., 2009); $^{44}{\rm Ti}$/Ti, $^{49}{\rm V}$/V and $^{51}{\rm Cr}$/Cr (Young et al., 1981; Webber, 1981; Leske, 1993; Connell, 2001b); $^{53}{\rm Mn}$/Mn (Young et al., 1981; Webber, 1981; Leske, 1993; DuVernois, 1997; Connell, 2001b), $^{55}{\rm Fe}$/Fe (Young et al., 1981; Wiedenbeck et al., 2001; Connell, 2001b), $^{56}{\rm Ni}$/Ni (Young et al., 1981), $^{57}{\rm Co}$/Co (Leske, 1993; Connell, 2001b), $^{59}{\rm Ni}$/Ni (Connell, 2001b); $^{67}{\rm Ga}$/Ga, $^{71}{\rm Ge}$/Ge and $^{73}{\rm As}$/As (Binns et al., 2022). The dispersion between the data point values is much larger than the weighted errors for $20\leq Z\leq 26$ isotopes, and $Z\geq 28$ isotopes have only one data point.
$Z$ Isotopic ratio Weighed mean (error)
4 $^{7}{\rm Be}$/Be 0.57 ($\pm$ 0.06)
18 $^{37}{\rm Ar}$/Ar 0.03 ($\pm$ 0.02)
20 $^{41}{\rm Ca}$/Ca 0.07 ($\pm$ 0.02)
22 $^{44}{\rm Ti}$/Ti 0.01 ($\pm$ 0.003)
23 $^{49}{\rm V}$/V 0.40 ($\pm$ 0.05)
24 $^{51}{\rm Cr}$/Cr 0.24 ($\pm$ 0.02)
25 $^{53}{\rm Mn}$/Mn 0.44 ($\pm$ 0.03)
26 $^{55}{\rm Fe}$/Fe 0.04 ($\pm$ 0.02)
27 $^{57}{\rm Co}$/Co 0.32 ($\pm$ 0.20)
28 $^{56}{\rm Ni}$/Ni 0.04 ($\pm$ 0.03)
28 $^{59}{\rm Ni}$/Ni 0.02 ($\pm$ 0.02)
31 $^{67}{\rm Ga}$/Ga 0.07 ($\pm$ 0.04)
32 $^{71}{\rm Ge}$/Ge 0.11 ($\pm$ 0.07)
33 $^{73}{\rm As}$/As 0.36 ($\pm$ 0.23)
4 Detectability of EC decay in current data
The last step of our analysis is to assess whether the impact of EC-decay is large enough to be seen in current CR data. This impact was calculated at IS energies and for the isotopic fluxes in the previous sections. However, the most precise CR data are elemental fluxes or ratios, in which the decayed fraction is diluted by the isotopic fraction (in the element) of the EC-unstable species. In the context of this simplified study, this limits the analysis to elemental fluxes for which the CR isotopic abundances of the EC-unstable nuclei have been measured, gathered in Table 5. Moreover, we also need to account for Solar modulation, as almost all data are taken in the Solar neighbourhood. We discuss below these various cases, based on the CR data extracted from the CRDB [^3] https://lpsc.in2p3.fr/crdb (Maurin et al., 2014, 2020, 2023) python library [^4] https://pypi.org/project/crdb/.
Table 3: Detectability of EC-decay (in the 1D-DM) with current IS CR data. The first three columns show the charge, name, and half-life of the EC-unstable species (for which IS elemental flux data exist). The next three columns show the element name, energy range, and best precision reached by current elemental IS fluxes from Voyager 1 (Cummings et al., 2016). We then show the expected decayed fraction at 20 MeV/n IS energy for the isotopic (next-to-last column) and the elemental fluxes (last column). The latter is calculated from the isotopic one, accounting for the isotopic fractions reported in Table 2. We stress that both 56Ni and 59Ni contribute to the Ni flux.
EC-unstable species Elemental IS data $f^{\rm IS}_{\rm dec}$ (20 MeV/n)
$Z$ Name $t_{1/2}$ Qty $E_{\rm k/n}$ [MeV/n] Prec. (in nuc) (in $Z$)
4 $^{7}{\rm Be}$ 53.22 d Be [61-96] 17% 3% 2%
18 $^{37}{\rm Ar}$ 35.011 d Ar [117-176] 22% 99% 3%
20 $^{41}{\rm Ca}$ 99.4 kyr Ca [18-193] 15% 27% 2%
22 $^{44}{\rm Ti}$ 59.1 yr Ti [75-199] 17% 95% 1%
23 $^{49}{\rm V}$ 330 d V [129-203] 22% 99% 40%
24 $^{51}{\rm Cr}$ 27.7015 d Cr [19-205] 17% 99% 24%
25 $^{53}{\rm Mn}$ 3.7 Myr Mn [153-209] 20% 14% 6%
26 $^{55}{\rm Fe}$ 2.7562 yr Fe [16-280] 10% 99% 4%
28 $^{56}{\rm Ni}$ 6.075 d Ni [21-221] 21% 100% 4%
$^{59}{\rm Ni}$ 81.82 kyr 21% 65% 1%
4.1 EC-decay in IS data
Only two human-made instruments ventured outside the Solar cavity, namely Voyager 1 (Stone et al., 2013) and Voyager 2 (Richardson et al., 2019), both launched in 1977. Cummings et al. (2016) published unmodulated elemental fluxes for H up to Ni (but not Co) from Voyager 1 in the so-called VLIS (very local interstellar medium). We show in Table 3 the corresponding EC-unstable isotopes (first 3 columns), along with the energy range (few tens to two hundreds MeV/n) and best precision ($\sim 20\%$) reached by the Voyager instruments on the associated fluxes (next 3 columns). The next-to-last column is our calculated decayed fraction, calculated at 20 MeV/n (the lowest Voyager data point). The latter is multiplied by the CR isotopic fractions (see Table 2) to obtain the expected decayed fraction in the elemental flux (last column). To be detectable, this fraction must be larger than the instrument precision: only V and Cr have significant decayed fractions, and V is the best candidate (to be precise, V first data point is at 129 MeV/n, for which the decayed fraction of 49V is 78% and that of V is 31%). On closer inspection, the $\sim 20\%$ CR data precision (on $Z\leq 26$ fluxes) automatically excludes all species with isotopic fractions smaller than this number. Among, the remaining ones, 49V and 51Cr are the most favourable (short-lived nuclei with $\sim 100\%$ decayed fractions), while 7Be is too light to attach electrons (and decay), and 53Mn is too long-lived to significantly decay. Note that 57Co ought to be a good candidate (short-lived with an isotopic fraction of 0.32), but Co is not available in Voyager data, unfortunately. We stress that the CR isotopic fraction data have a dispersion (not shown) much larger than the weighted error reported in Table 2, which could change the detectability of EC-decay in IS V and Cr fluxes.
Table 4: Energy range and best achieved precision on existing CR TOA for isotopic measurements (first three columns), either for EC-unstable parent-to-daughter ratios (Scott, 2005), or for isotopic ratios (same references as in Table 2).
EC-unstable species Parent/daughter (TOA data) Isotopic ratio (TOA data)
$Z$ Name $t_{1/2}$ Qty $E_{\rm k/n}$ [GeV/n] Prec. Qty $E_{\rm k/n}$ [GeV/n] Prec.
4 $^{7}{\rm Be}$ 53.22 d $^{7}{\rm Be}$$^{7}{\rm Li} - -$ $^{7}{\rm Be}$/Be [0.08-1.42] 2%
18 $^{37}{\rm Ar}$ 35.011 d $^{37}{\rm Ar}$$^{37}{\rm Cl}$ [0.12-0.35] 9% $^{37}{\rm Ar}$/Ar [0.49-0.57] 6%
20 $^{41}{\rm Ca}$ 99.4 kyr $^{41}{\rm Ca}$$^{41}{\rm K}$ [0.12-0.37] 10% $^{41}{\rm Ca}$/Ca [0.58-0.65] 6%
22 $^{44}{\rm Ti}$ 59.1 yr $^{44}{\rm Ti}$$^{44}{\rm Ca}$ [0.13-0.39] 14% $^{44}{\rm Ti}$/Ti [0.32-0.69] 27%
23 $^{49}{\rm V}$ 330 d $^{49}{\rm V}$$^{49}{\rm Ti}$ [0.14-0.41] 8% $^{49}{\rm V}$/V [0.32-0.69] 19%
24 $^{51}{\rm Cr}$ 27.7015 d $^{51}{\rm Cr}$$^{51}{\rm V}$ [0.14-0.42] 7% $^{51}{\rm Cr}$/Cr [0.32-0.72] 10%
25 $^{53}{\rm Mn}$ 3.7 Myr $^{53}{\rm Mn}$$^{53}{\rm Cr}$ [0.14-0.42] 9% $^{53}{\rm Mn}$/Mn [0.32-0.74] 10%
26 $^{55}{\rm Fe}$ 2.7562 yr $^{55}{\rm Fe}$$^{55}{\rm Mn}$ [0.14-0.43] 11% $^{55}{\rm Fe}$/Fe [0.62-0.94] 10%
28 $^{56}{\rm Ni}$ 6.075 d $^{56}{\rm Ni}$$^{56}{\rm Fe} - -$ $^{56}{\rm Ni}$/Ni [0.65-0.99] 74%
27 $^{57}{\rm Co}$ 271.811 d $^{57}{\rm Co}$$^{57}{\rm Fe}$ [0.15-0.44] 16% $^{57}{\rm Co}$/Co [0.32-0.32] 30%
28 $^{59}{\rm Ni}$ 81.82 kyr $^{59}{\rm Ni}$$^{59}{\rm Co} - -$ $^{59}{\rm Ni}$/Ni [0.10-0.50] 70%
31 $^{67}{\rm Ga}$ 3.2617 d $^{67}{\rm Ga}$$^{67}{\rm Zn} - -$ $^{67}{\rm Ga}$/Ga [0.13-0.70] 58%
32 $^{71}{\rm Ge}$ 11.43 d $^{71}{\rm Ge}$$^{71}{\rm Ga} - -$ $^{71}{\rm Ge}$/Ge [0.13-0.70] 64%
33 $^{73}{\rm As}$ 80.30 d $^{73}{\rm As}$$^{73}{\rm Ge} - -$ $^{73}{\rm As}$/As [0.13-0.70] 65%
4.2 From IS to TOA decaying fractions
All previous calculations were performed on IS energies. When we consider TOA fluxes, the generic source term no longer factors out in the decayed fraction. Indeed, Solar modulated fluxes critically depend on the spectral shape of the IS fluxes. We can go around this issue as follows.
  1. 1.
    We assume that the variability of any IS flux shape is encompassed by the variability of He to Ni IS fluxes, as parametrised in Shen et al. (2019, 2025). We denote the corresponding fluxes by $\psi^{\rm IS}_{\rm Chen}$.
  2. 2.
    We model the decayed fraction by a generic sigmoid function
    $\Sigma=(1-f_{\rm min})\times\left[1-\frac{1}{1+\exp\left[-k\log(E_{\rm k/n})-\log(E_{0}))\right]}\right]\,.$
    We checked that the latter reproduces the variety of decayed fraction shapes. The parameters of the sigmoid control the maximal decayed fraction at low energy, $f_{\rm min}$, and the width, rise, and energy position at which the sigmoid goes to zero at a few GeV/n.
  3. 3.
    We calculate the TOA decaying fraction by modulating $\psi^{\rm IS}_{\rm no\penalty 10000\ decay}=\psi^{\rm IS}_{\rm Chen}$ and $\psi^{\rm IS}_{\rm EC\penalty 10000\ decay}=(1-\Sigma)\,\psi^{\rm IS}_{\rm no\penalty 10000\ decay}$ (TOA flux without and with decay, respectively), and then forming their relative difference; this gives $f^{\rm TOA}$.
We checked (not shown) that $f^{\rm TOA}$ is insensitive to the IS spectral shape (i.e., whether it is a primary or secondary species) and to the sigmoid parameters (i.e., whether it has a mild, strong, steep or flat decrease). All together, the TOA decayed fraction matches the IS one with the energy shifted by the Solar modulation parameter (see Eq. 4):
$f^{\rm TOA}(E_{\rm k/n}^{\rm TOA})=f^{\rm IS}(E_{\rm k/n}^{\rm IS}=E_{\rm k/n}^{\rm TOA}+\phi_{\rm FF}\cdot Z/A)\;.$
The above relation implies that, for the same energy $f^{\rm TOA}\!<\!f^{\rm IS}$. It also means that $f^{\rm TOA}$ is larger for periods of low Solar activity than for high Solar acticity ones).
4.3 EC-decay in TOA isotopic data
TOA CR data have been obtained for different levels of Solar activity. In this section, for simplicity and to be conservative, we model the decayed fraction for a low level of Solar activity, $\phi_{\rm FF}=500$ MV, corresponding to an energy shift of $\sim 250$ MeV/n, and thus a TOA decayed fraction as large as possible.
Data status
As already said, isotopic separation is hard to achieve in CR experiments. Furthermore, fluxes are more difficult to measure than ratios: the former requires an excellent knowledge of the detector acceptance, whereas the latter does not (and also cancels out many systematics). As a result, almost all isotopic CR data are ratios, that we gather in Table 4. First, we have isotopic ratios (4th to 6th columns), formed from the EC-unstable parent to their daughter isotope. These ratios maximise the decay impact (decayed parent in the numerator and EC-fed daughter in the denominator), but are not always provided by the experiments. The largest and most precise dataset is that of ACE-CRIS (Scott, 2005), with a precision varying from 7% to 16%. Second, we also have isotopic ratios (last 3 columns), which are more systematically provided. We clearly see the pattern of decreasing precision with the isotope mass, as both their abundance and isotopic separation ($\Delta A/A$) decrease with $A$, making them harder to measure [^5] Actually, the data precision is a non-trivial balance between the instrument mass resolution, the isotopic abundance of the EC-unstable species, and the relative abundance of the latter with respect to the neighbour isotopes. The trend seems to be dominated by the mass resolution and abundance of the EC-unstable under scrutiny, with a poorer precision for heavier isotopes.. These ratios have been measured by numerous experiments, the most precise one being ACE-CRIS’ (Connell, 2001b; Ogliore et al., 2009; Binns et al., 2022).
Figure 6: Decaying fraction of EC-unstable isotopes (lines) at $\phi_{\rm FF}=500$ MV as a function of TOA energies, compared to the TOA CR data precision (rectangular boxes). We do not show 7Be, 41Ca, 53Mn, and 59Ni, whose decayed fraction is below the percent level, and thus largely below the data precision. In order not to overcrowd the figure, we also grouped species whose CR data are of similar precision (from Table 2) in the same rectangles.
EC-decay detectability
We show in Fig. 6 the comparison between the decayed fraction (lines) and the best data precision (filled rectangles). As before, we say that EC-decay is detectable if the decaying fraction of an isotope is larger than its data relative precision. First, light EC-unstable species are very rare, with only $^{7}{\rm Be}$ for $Z<15$. Even though the isotopic ratio of the latter is measured with the greatest precision ($\sim 2\%$), its decay fraction is less than 1% (owing to its long attachment time). Second, for $Z>15$, we have different behaviours depending on the isotopes half-life. The decay fraction for intermediate-lived isotopes is strongly dependent on the decay time: 41Ca ($t_{1/2}=99.4$ kyr), 53Mn ($t_{1/2}=3.7$ Myr) and 59Ni ($t_{1/2}=81.82$ kyr) have a decay fraction at the percent level (not shown) above a few hundreds of MeV/n, while 44Ti ($t_{1/2}=59.1$ yr) reaches 10% (solid cyan line). For short-lived isotopes ($t_{1/2}\lesssim$ yr), the decaying fraction strongly depends on the atomic number, with the heaviest species (67Ga, 71Ge and 73As) showing the highest decayed fraction.
Table 5: Minimal energy and best precision achieved on elemental CR TOA data, either for the fluxes ($4^{\rm th}$ to $6^{\rm th}$ columns) or the parent-to-daughter elemental flux ratios (last 3 columns). For $Z\leq 30$, the most precise data are from ACE-CRIS (Lave et al., 2013) and AMS-02 (Aguilar et al., 2021a, b); note that AMS-02 data for $22<Z<25$ and $Z>26$ have not been released yet. For Ga/Zn, Ge/Ga and As/Ge, only a few data points exist from TIGER (Rauch et al., 2009), Super-TIGER (Murphy et al., 2016), ACE-CRIS (Binns et al., 2022) and CALET (Adriani et al., 2025). More CR data exist for heavier elemental ratios, but their isotopic content is not measured, so that the decayed fraction cannot be estimated. Furthermore, we stress that the best precision reported in this table is not necessarily obtained at the lowest energy reported.
EC-unstable species Elemental flux (TOA data) Elemental ratio (TOA data)
$Z$ Name $t_{1/2}$ Qty $E_{\rm k/n}$ [GeV/n] rel.err Qty $E_{\rm k/n}$ [GeV/n] rel.err
4 $^{7}{\rm Be}$ 53.22 d Be 0.06 2% Be/Li 0.08 6%
18 $^{37}{\rm Ar}$ 35.011 d Ar 0.12 3% Ar/Cl 0.12 6%
20 $^{41}{\rm Ca}$ 99.4 kyr Ca 0.14 3% Ca/K 0.14 5%
22 $^{44}{\rm Ti}$ 59.1 yr Ti 0.14 3% Ti/Ca 0.14 5%
23 $^{49}{\rm V}$ 330 d V 0.14 4% V/Ti 0.14 6%
24 $^{51}{\rm Cr}$ 27.7015 d Cr 0.14 3% Cr/V 0.14 6%
25 $^{53}{\rm Mn}$ 3.7 Myr Mn 0.14 4% Mn/Cr 0.14 5%
26 $^{55}{\rm Fe}$ 2.7562 yr Fe 0.17 2% Fe/Mn 0.17 5%
28 $^{56}{\rm Ni}$ 6.075 d Ni 0.17 5% Ni/Fe 0.17 5%
27 $^{57}{\rm Co}$ 271.811 d Co 0.17 5% Co/Fe 0.17 5%
28 $^{59}{\rm Ni}$ 81.82 kyr Ni 0.17 5% Ni/Co 0.17 5%
31 $^{67}{\rm Ga}$ 3.2617 d Ga - - Ga/Zn 0.30 12%
32 $^{71}{\rm Ge}$ 11.43 d Ge - - Ge/Ga 0.30 13%
33 $^{73}{\rm As}$ 80.30 d As - - As/Ge 0.30 21%
All in all, the detectability of EC-decay is possible if the curves are below the corresponding boxes (of same colour). This only clearly happens for 51Cr and 55Fe (dashed and dotted red lines), and otherwise, we barely meet the condition for 57Co (thin cyan line), and 67Ga, 71Ge and 73As (dashed, dotted, and dash-dotted blue lines). So it seems that we are at the verge of being able to detect EC-decay in CR data. However, this must be mitigated by the fact that EC-decay is overestimated in our modelling, as we neglected energy losses. And this is without considering the impact of nuclear production (Clinton and Waddington, 1993; Waddington, 1996, 1998) and inelastic cross-section uncertainties (Maurin et al., 2025), possibly larger than the decayed fraction itself.
4.4 EC decay in TOA elemental data
It is also interesting to study our results in the context of available data on elemental fluxes, which are generally easier to measure (they bypass the challenges of isotopic separation), and therefore can be obtained with higher precision.
Data status
Elemental fluxes have been measured by a huge number of experiments. But at low energy and for $Z\leq 30$, the most precise data have a few percent precision, and are obtained down to $\simeq 100$ MeV/n from ACE-CRIS (Lave et al., 2013) and AMS-02666To be precise, AMS-02 data are in rigidity unit, and converting them in kinetic energy per nucleon introduces additional systematics at the few percent level (Derome et al., 2019), because the isotopic content is not measured above a few GeV/n. with a minimum energy of $\simeq 500$ MeV/n (Aguilar et al., 2021a, b). A summary view of the relevant data are gathered in Table 5, for the elemental fluxes (4th to 6th columns), but also for ratios of parent/daughter elemental fluxes (last 3 columns) – the latter ratios contain the EC-unstable isotope in their numerator and its daughter in their denominator, maximising the EC-decay impact. Only ratios have been measured for $Z>30$, by TIGER (Rauch et al., 2009), Super-TIGER (Murphy et al., 2016), ACE-CRIS (Binns et al., 2022) and CALET (Adriani et al., 2025). These experiments provide data up to $Z\simeq 40$, but that cannot be exploited here: the isotopic content of EC-unstable species is necessary to obtain the elemental decaying fraction, and this content has only been measured [^7] In principle, we could use the predicted isotopic content instead, but we do not have it in our very simplified approach. for Z=3133Z=31-33 (Binns et al., 2022).
EC-decay detectability
Figure 7: Same as in Fig. 6, but for the decayed fraction in TOA elemental fluxes (compared to TOA CR data). We do not show Be, Ca, Ti, Mn, and Ni, whose decayed fraction is below the sub-percent level, and thus largely below the data precision.
As discussed in the previous section, the TOA decayed fraction of intermediate-lives isotopes 7Be, 41Ca, 53Mn, and 59Ni are at the percent level, so when diluted by their isotopic content (0.57, 0.07, 0.44 and 0.02, respectively, see Table 2), this fraction is sub-percent in the elemental TOA fluxes (Be, Ca, Mn, and Ni, respectively). Even the decayed fraction in Ti is sub-percent, owing to its small isotopic fraction: the decayed fraction of 44Ti is 10%, diluted by its 0.01 isotopic fraction. Only short-lived EC-unstable isotopes provide $\gtrsim 1\%$ decayed fractions. We show, in Fig. 7, their expected decayed fractions (lines), compared to the CR data precision (rectangular boxes). For $Z\leq 30$, only V, Cr, and Co (thin-dashed, dotted, and thick-dashed blue lines, respectively) have data precision at the level or better than the expected decayed fraction, Co being the most favourable case. Actually, forthcoming AMS-02 data on these species will start at $\gtrsim 500$ MeV/n with a $\gtrsim 6\%$ (uncertainties dominated by systematics on the acceptance), so ACE-CRIS data are more favourable to search for this decay. For $Z>30$, neither Ga and Ge (red lines), nor As (cyan line) meet this requirement, and by a factor of two or more.
5 Conclusions
We have revisited the impact of EC decay on isotopic and elemental fluxes, assessing the detectability of this decay in current IS and TOA CR data.
First, on the modelling side, we obtained analytical solutions accounting for multiple ionic states of CRs during propagation. We showed that the widely used 2-level approximation (i.e., fully ionised CR or single electron attached) differs from the full calculation (by $\lesssim 50\%$) only for intermediate-lived nuclei $t_{1/2}\sim{\cal O}(10\,{\rm kyr})$ with $Z>40$ and at IS energies $\lesssim 100$ MeV/n. There are actually very few EC-unstable species in this decay time and charge range. Furthermore, intermediate-lived nuclei are found to have mild decaying fractions, well below the precision of current (and possibly forthcoming) data. As a result, for all practical purpose, the 2-level approximation is deemed appropriate.
Second, we have made a quantitative study of the impact of attachment and stripping cross-section uncertainties on the decaying fractions. Short-lived species (i.e., $t_{1/2}<{\cal O}({\rm kyr})$) are very sensitive to uncertainties on the attachment cross-sections (timescales $\gtrsim$ kyr), but insensitive to those on stripping cross-sections. The latter only impacts intermediate-lived nuclei (also impacted by attachment cross-section uncertainties), but only mildly. We stress that the cross-sections used in the literature and in this paper were derived more than forty years ago. While their values are not critical, unless they are off by a factor of a few, updating them and assessing their accuracy and precision would certainly be an improvement.
Third, we have compared the decaying fraction in the still widely used LBM (for $Z>30$ analyses) and in a more realistic diffusion model. The leakage-lifetime approximation (at the heart of the LBM) is known to fail for $\beta$-unstable species, and it also fails for EC-unstable species. Nevertheless, the decaying fraction trend as a function of energy, charge, and decay times, are qualitatively similar in both models. Differences on the fluxes can be as high as a factor of a few, but in a parameter space where very few EC-unstable species actually lie.
Last, we have compared the decayed fraction to both IS and TOA current data precision. While IS data are more favourable in terms of energy, only elemental fluxes – in which the isotope decay is diluted by the presence of the other stable isotopes – have been measured. Low-energy TOA data are more numerous and benefit from the latest generation of high-precision experiments (mainly ACE-CRIS for isotopes and fluxes, and AMS-02 for fluxes). All in all, only 49V, 51Cr and 57Co meet the EC-decay detectability condition – i.e., the predicted decayed fraction is larger than the current data precision –, in both isotopic TOA data and elemental IS or TOA fluxes. We stress that all our calculations were performed neglecting energy losses, so that the decayed fractions are probably slightly overestimated.
In summary, while the impact of EC decay remains difficult to confirm in current CR, the isotopes of V, Cr, and Co are the most promising candidates to do so. Several aspects of our calculation require refinement, though: first, the model must incorporate energy losses and solve the transport equation with the complete network of nuclear production cross-sections; second, uncertainties on nuclear cross-sections for $Z>30$ must be quantified, as they are known to bring large uncertainties in the modelling, especially for pure secondary species. In any case, the presence of many short-lived EC-unstable isotopes (with $Z>30$) that should be fully decayed, combined with the prospect of measurements in the near future (e.g., from ISS-TIGER), provides a strong motivation to investigate further this topic.
Acknowledgments
This work is part of the project “Cosmic ray antideuterons as a probe for new physics” with project number OCENW.KLEIN.387 (Budget Number 11680) of the research programme Grant Open Competition Domain Science, which is financed by the Dutch Research Council (NWO).
Appendix A Attachment and stripping cross-sections
Electron attachment and stripping cross-sections are essential to properly describe the transport of EC-unstable species. We start below with the formulae for a single electron attachment (hydrogen-like ions with only two possible charged states) and discuss the corrections needed to extend them to a multilevel model. We recall that these cross-sections are summed over the relevant ISM component densities, see Eq. (5).
Electron attachment is denoted electron capture in the literature (Eichler and Stöhlker, 2007), but to avoid confusion with the EC-decay process, we keep the former denomination throughout this paper. We use (and report below) the cross-section parametrisations of Letaw et al. (1984b), derived in the context of CR studies. The value and precision of these cross-sections should probably be revisited, in the light of the progress made in EC in ion-atom collisions; see Tolstikhina and Shevelko (2023) for a recent, comprehensive, though broader review. But here, we keep these parametrisations for comparison purpose, as they have been used in all previous CR studies. We nevertheless provide a quantitative estimate of the impact (on the propagated CR fluxes) of these cross-section uncertainties in 3.3, to assess how critical they are.
A.1 Attachment
The capture of an electron, from a neutral ISM target T, by a CR projectile P, is a combination of radiative attachment (ra) and non-radiative attachment (nra),
$\sigma^{\rm att}_{\rm P+T}=\sigma^{\rm ra}_{\rm P+T}+\sigma^{\rm nra}_{\rm P+T}\,$, (10)
with
${\rm P}^{q+}+{\rm T}$ $\rightarrow$ ${\rm P}^{(q-1)+}+{\rm T}^{+}+\gamma\,\quad{\rm(ra)}\,$,
${\rm P}^{q+}+{\rm T}$ $\rightarrow$ ${\rm P}^{(q-1)+}+{\rm T}^{+}\quad\quad\;\,{\rm(nra)}\,.$
In the above reactions, $q+$ is the initial charge of the projectile, with $q+=Z_{\rm P}$ if the projectile is completely ionised. In the non-radiative process, the excess momentum of the electron is carried out by the photon (3-body problem). In the non-radiative one, the momentum is conserved through the recoil of the target T (from which the electron was captured). These two processes have similar energy-dependences in the asymptotic relativistic regime, with $\sigma^{\rm ra}_{\rm P+T}\propto Z_{\rm P}^{5}Z_{\rm T}/E$ and $\sigma^{\rm nra}_{\rm P+T}\propto Z_{\rm P}^{5}Z_{\rm T}^{5}/E$, but strongly differ in the highly non-relativistic one, with $\sigma^{\rm ra}_{\rm P+T}\propto\beta^{-5}$ and $\sigma^{\rm nra}_{\rm P+T}\propto\beta^{-12}$ – for a thorough discussion, see Eichler and Stöhlker (2007)
Extension of the 1s–1s attachment formula
The starting point is to consider the attachment of an electron from the K-shell (or subshell 1s) of a single-electron target atom, to the K-shell of a fully stripped projectile, $\sigma^{\rm att}_{\rm P(1s)\,+\,T(1s)}$, whose formula is detailed in the next paragraph. Accounting for the attachment in all other projectile shells – which falls off as $1/n^{3}$ for the $n$-th shell (Crawford, 1979) –, we have
$\sigma^{\rm att}_{\rm P\,+\,T(1s)}=\sum_{n=1}^{\infty}\frac{1}{n^{3}}\,\sigma^{\rm att\;(1s)}_{\rm P(1s)\,+\,T(1s)}=1.202\,\sigma^{\rm att\;(1s)}_{\rm P(1s)\,+\,T(1s)}\,.$ (11)
In the case of targets with more than one K-shell electron (or electrons also in other shells), one has to take into account the extra number of electrons available, and also the partial screening of the nuclear potential due to the presence of these extra electrons. This screening leads to an effective charge of the target (Clementi and Raimondi, 1963; Clementi et al., 1967; Guerra et al., 2017), with $Z_{\rm T}^{\rm eff}\approx Z_{\rm T}-0.3$ a reasonable approximation for practical calculations. In the context of H and He targets in the ISM, the relevant formulae to use are
$\sigma^{\rm att}_{\rm P+HI}$ $=$ $1.202\,\sigma^{\rm att}_{\rm P(1s)\,+\,T(1s)}(Z_{\rm T}=1)\,;$ (12)
$\sigma^{\rm att}_{\rm P+He}$ $=$ $2\times 1.202\,\sigma^{\rm att}_{\rm P(1s)\,+\,T(1s)}(Z_{\rm T}^{\rm eff}=1.7)\,.$ (13)
1s–1s radiative attachment
The cross-section $\sigma^{\rm ra}_{\rm P(1s)\,+\,T(1s)}$ can be calculated from the photoionisation cross-section (Crawford, 1979),
${\rm P}^{q+}+\gamma\rightarrow{\rm P}^{(q+1)+}+e^{-}\,$,
using the method of detailed balance, assuming there are two electrons available in the K-shell. As a consequence, this cross-section has to be divided by two to get the one for the attachment of a single electron. For a fully stripped projectile, this cross-section reads (Crawford, 1979; Letaw et al., 1984b):
$\sigma^{\rm ra}_{\rm P(1s)\,+\,T(1s)}\!\!$ $=$ $\!\!\frac{3}{2}\,\sigma_{T}\,Z_{\rm P}\,Z_{\rm T}\,\frac{\beta\,\gamma}{(\gamma-1)^{3}}\,A^{4+2\xi}\,\exp\left(\frac{-2A}{\beta\cos A}\right)$ (14)
$\times$ $\bigg(M(\beta)\;\Big[1\,+\,R(A)\Big]\,+\pi\,A\,N(\beta)\bigg)\,$,
where $\sigma_{T}$ is the Thompson cross-section, $Z_{\rm P}$ is the projectile charge, $Z_{\rm T}$ the target charge, $A=\alpha\,Z_{\rm P}$ (with $\alpha=0.00729$ the fine structure constant), and
$\xi$ $=$ $\left(1\,-\,A^{2}\right)^{1/2}-1\,$,
$M(\beta)$ $=$ $\frac{4}{3}+\frac{\gamma\,(\gamma-2)}{(\gamma+1)}\,\left[1-\frac{1}{2\beta\gamma^{2}}\,\ln\left(\frac{1+\beta}{1-\beta}\right)\right]\,$,
$N(\beta)$ $=$ $\frac{1}{15\beta^{3}}\left(-4\gamma+34-\frac{63}{\gamma}+\frac{25}{\gamma^{2}}+\frac{8}{\gamma^{3}}\right)$
$-\frac{(\gamma-2)}{2\beta^{2}\gamma(\gamma\!+\!1)}\ln\left(\frac{1\!+\!\beta}{1\!-\!\beta}\right)\,$,
$R(A)$ $=$ $-\exp\left(-8.4\,A^{2}+14\,A-8.28\right)\,.$
1s–1s non-radiative attachment
A non-relativistic quantum mechanical treatment (OBK approximation) allows for computing the probability of transferring an electron from the K-shell of a single-electron target atom to the K-shell of a fully stripped nucleus (Wilson, 1978):
$\sigma^{\rm nra\;(1s-1s)}_{\rm P+T}=\frac{\pi}{5}\,\frac{2^{18}\,a_{0}^{2}\,(Z_{\rm P}\,Z_{\rm T})^{5}\,\gamma^{2}\,S^{8}}{\left[S^{2}\!+\!(Z_{\rm P}+Z_{\rm T})^{2}\right]^{5}\,\left[S^{2}+(Z_{\rm P}-Z_{\rm T})^{2}\right]^{5}}\,$, (15)
where $S=\beta\gamma/\alpha$ and $a_{0}=5.29\times 10^{-9}$ cm is the Bohr radius.
A.2 Stripping
Stripping processes can be schematically represented as follows:
${\rm P}^{q+}+{\rm T}$ $\rightarrow$ ${\rm P}^{(q+1)+}+{\rm T}+e^{-}\,.$
${\rm P}^{q+}+{\rm T}^{p+}$ $\rightarrow$ ${\rm P}^{(q+1)+}+{\rm T}^{(p-1)+}\,.$
The cross-section for this second process is almost negligible (Crawford, 1979), and the first one is calculated from Mott and Massey’s relativistic ionisation cross-section (Wilson, 1978). The loss of a K-shell electron by a hydrogenic atom is given by
$\sigma^{\rm str}_{\rm P+T}=4\,\pi a_{0}^{2}\,\left(\frac{\alpha}{Z_{\rm P}\beta}\right)^{2}\,(Z_{\rm T}^{2}+Z_{\rm T})\,C_{1}\,\left(\ln\frac{4\beta^{2}\gamma^{2}}{C_{2}\alpha^{2}Z_{\rm P}^{2}}-\beta^{2}\right)$ (16)
where $C_{1}=0.285$ and $C_{2}=0.048$ are constants for capture from K-shell. The $Z_{\rm T}$ term arises from a sum over the contributions from all individual electrons in the target atom, while the $Z_{\rm T}^{2}$ term accounts for the ionisation by the nucleus. The final stripping cross-section formula is the weighted sum of the cross-sections for the different ISM target abundances.
A.3 Extension for the multi-level model
All the formulas presented in the previous sections consider just two ionised states for the projectile, with effective charge $q_{\rm P}=Z_{\rm P}$ and $q_{\rm P}=Z_{\rm P}-1$. Indeed, they were specifically derived for attachment of an electron to a fully ionised nucleus or stripping from a hydrogenic atom. In our analysis, we generalise the transport equation to a multi-level model (see B.1), tracking the evolution of the CR ions up to any level of ionisation state $q_{\rm P}\in\{Z_{\rm P},(Z_{\rm P}-1),(Z_{\rm P}-2),\dots\}$. This requires the knowledge of the attachment and stripping cross-sections for partly ionised ions. The semi-empirical Schlachter formula proposes $\sigma^{\rm att}\propto q_{\rm P}^{3.9}$ (Schlachter et al., 1983), with an accuracy $\lesssim 2$ (Tolstikhina and Shevelko, 2023). As there are large uncertainties on this formula, we keep the description simple and assume the above formulae apply for all ionisation states. This assumption was actually made in the context of the more complex problem of the study of charge states of relativistic heavy ions in matter (Scheidenberger et al., 1998). In any case, we find that only attachment levels up to a few are relevant for most of the species in the CR context (see Sect. 3.4), so this approximation should not be too critical. Another assumption we make is to neglect multielectron capture or stripping – we assume electrons are attached or stripped one at a time –, whereas it can be significant at almost all energies (Tolstikhina and Shevelko, 2023). However, the dependence of the multielectron cross-sections on several parameters (atomic structure of colliding particles, velocity of the incident ion, and other atomic parameters) is still being investigated experimentally and theoretically (e.g., Wu et al., 2025). In the end, for the same reasons as above – simplicity and the fact that mostly low-level attachments and stripping seem to be relevant for CR –, this assumption should be good enough in the CR context.
Appendix B Solutions of the transport equation
The transport equations, for a $\beta$-unstable species and a generic source term $Q$, in the LBM and thin-disc 1D-DM, are given respectively by
$\frac{N}{\tau_{\rm esc}}+\left(\Gamma^{\rm inel}+\Gamma^{\beta}\right)N=Q\,$, (17)
and
$D\frac{d^{2}N}{dz^{2}}+\left(2h\delta(z)\,\Gamma^{\rm inel}+\Gamma^{\beta}\right)N=2h\delta(z)\,Q\,.$ (18)
The (semi-) analytical solutions of the transport equation, for stable and $\beta$-radioactive GCR ions, (with) without energy redistribution, are gathered in Maurin (2020), for the LBM and 1D-DM. In this appendix, we extend these solutions to the EC case, without energy redistribution and without galactic wind (for the 1D-DM).
Compared to the standard solutions, we now have to follow the various ionisation states of the GCR species. Indeed, whereas CRs are expected to be fully ionised at high-enough energy (see Fig. 1 in Sect. 2.3), several ionisation states may co-exist at lower energy (Letaw et al., 1984b, 1985). As a result, instead of solving the transport equation for the differential density of the CR ion $N$, we have to solve it for a tower of ionised states, denoted below $N_{0}$ (fully ionised), $N_{1}$ (one $e^{-}$ attached), $N_{2}$ (two $e^{-}$ attached), etc. Each level is connected to the one below and above via the electron attachment and stripping rates, $\Gamma^{\rm att}$and $\Gamma^{\rm str}$, the decay rate $\Gamma^{\rm EC}=1/(\gamma\tau_{EC})$ being enabled for all $N_{j>0}$ (i.e., at least one electron attached). The total measured flux of this CR species is then given by
$N_{(n+1)\penalty 10000\ {\rm lev}}=N_{0}+N_{1}+N_{2}+\dots+N_{n}\,$, (19)
where $n\leq Z$ (with $Z$ the charge of the CR under scrutiny) is the enforced truncation level in the calculation.
B.1 Multi-level solution for EC decay
We consider below a single EC-unstable CR species, and we assume a generic source term $\cal S$} (of primary or secondary origin), providing the ionised state $N_{0}$ only. In this first step, we neglect energy redistribution terms, and for simplicity, we also discard the convective term in the 1D-DM. The multi-level (up to level $n$) transport equation for the coupled ionised states, for any of the above models, can be written generically (with $j=1,\dots,n-1$)
[…]
$\left({\cal O},\,{\cal F},\,{\cal S}\right)_{\rm LBM\;}$ $=\left(\tau_{\rm esc}^{-1},\;1,\;Q\right)\,$, (21)
$\left({\cal O},\,{\cal F},\,{\cal S}\right)_{\rm 1D-DM}$ $=\left(-D\,\partial^{2}_{z},\;2h\,\delta(z),\;2h\,\delta(z)\,Q\right)\,.$
The main differences between the 1D-DM and the LBM are the differential transport operator (instead of the escape time) and the attachment, stripping, inelastic cross-sections, and source term tied to the thin disc only instead of the full LB volume. Alternatively, EC decay is possible in the whole confinement volume in both cases.
For the last level at which the multi-level calculation stops, attachment is not enabled, in order to conserve the total number of CRs. Also, we have
$\tilde{\Gamma}^{\rm EC}=\begin{cases}\Gamma^{\rm EC}/2\quad{\rm for}\penalty 10000\ i=1\,,\\ \Gamma^{\rm EC}\quad{\rm otherwise}.\end{cases}$ (22)
Indeed, the EC half-life measured in the laboratory is for nuclei with two electrons in the K-shell, so that when only one electron is attached (i.e., $i=1$ above), the half-life is doubled and the rate halved. In the above equations, the energy and species (i.e., $A$ and $Z$) dependences are implicit, but are in almost all terms. Following A, we have $\Gamma^{\rm att}=\tilde{\Gamma}^{\rm ra}+\Gamma^{\rm nra}$ with
$\tilde{\Gamma}^{\rm ra}=\begin{cases}\Gamma^{\rm ra}\quad{\rm for}\penalty 10000\ n=1\,,\\ 2\Gamma^{\rm ra}\quad{\rm otherwise}.\end{cases}$ (23)
For the LBM, the above set of equations is directly a coupled set of algebraic equations. For the 1D-DM, this is a set of coupled differential equations, but following the procedure highlighted, e.g., in Maurin et al. (2001) – i.e., solving in the halo with boundary condition for each level, $N_{j}(z=L)=0$, and integrating through the thin disc ensuring continuity –, we also arrive at a similar set of coupled algebraic equations. The latter can be written in a matrix form, $\underline{\underline{A}}\;\underline{N}\,=\,\underline{S}$,
$\begin{bmatrix}\alpha_{0}&-\Gamma^{\rm str}&0&\cdots&0\\ -\Gamma^{\rm att}&\alpha_{1}&-\Gamma^{\rm str}&&\vdots\\ 0&\ddots&\ddots&\ddots&0\\[2.84526pt] \vdots&&-\Gamma^{\rm att}&\!\alpha_{n-1}\!&-\Gamma^{\rm str}\\[5.69054pt] 0&\cdots&0&-\Gamma^{\rm att}&\alpha_{n}\end{bmatrix}\;\begin{bmatrix}N_{0}\\[5.69054pt] N_{1}\\[5.69054pt] \vdots\\[5.69054pt] N_{n-1}\\[5.69054pt] N_{n}\end{bmatrix}=\begin{bmatrix}Q\\[5.69054pt] 0\\[5.69054pt] \vdots\\[5.69054pt] 0\\[5.69054pt] 0\end{bmatrix}\,$, (24)
where the $\alpha$ coefficients are defined to be
$\begin{cases}\alpha_{0}={\cal P}+\Gamma^{\rm inel}+\Gamma^{\rm att},\\ \alpha_{j}={\cal V}+\Gamma^{\rm inel}+\Gamma^{\rm str}+\Gamma^{\rm att}\quad(j=1,\dots,n-1),\\ \alpha_{n}={\cal V}+\Gamma^{\rm inel}+\Gamma^{\rm str},\end{cases}$ (25)
and with the multiplicative factors $\cal P$} and $\cal V$}, respectively given by (for the LBM and the 1D-DM)
$\left({\cal P},\;{\cal V}\right)_{\rm LBM\;}$ $=\left(\tau_{\rm esc}^{-1},\;\tau_{\rm esc}^{-1}+\tilde{\Gamma}^{\rm EC}\right)\,$, (26)
$\left({\cal P},\;{\cal V}\right)_{\rm 1D-DM}$ $=\left(\frac{D}{hL},\;\sqrt{\frac{D\,\tilde{\Gamma}^{\rm EC}}{h^{2}}}\coth\left(\sqrt{\frac{L^{2}\tilde{\Gamma}^{\rm EC}}{D}}\right)\right)\,.$
Formally, the solution of Eq. (24) is $\underline{N}\,=\,\underline{\underline{A}}^{-1}\,\underline{S}$. Owing to the simple form of $\underline{S}$, we have $N_{i}=Q\,(\underline{\underline{A}}^{-1})_{i,0}$. The coefficients of the inverse of a generic tridiagonal matrix can be found, for instance, in Usmani (1994) or Huang and McColl (1997). Defining recursively
$\theta_{j}=\alpha_{j}\,\theta_{j+1}-\Gamma^{\rm att}\Gamma^{\rm str}\,\theta_{j+2}\quad\quad(j=n-1,n-2,\ldots,0)$ (27)
[…]
$N_{j}=q\,\left(\prod\Gamma^{\rm att}_{1}\dots\Gamma^{\rm att}_{j-1}\right)\frac{\theta_{j+1}}{\theta_{0}}\quad(j=0,\dots,n)\,$, (28)
or alternatively
[…]
We stress that the dimension of the latter quantity is a power-law of time T, $\theta_{j}]=$ Tj-1-n, so that we recover $N_{j}]=[Q]$ T-1.
To summarise, the solutions can be computed recursively from Eqs. (28) and (27), with $\alpha_{i}$ defined in Eq. (25) from the multiplicative factors $\cal P$} and $\cal V$} defined in Eq. (26). This solution can be extended to a mixed unstable species, with both EC and $\beta$ decay, by replacing $\tilde{\Gamma}^{\rm EC}$ by $(\tilde{\Gamma}^{\rm EC}+\gamma^{\beta})$ in the $\cal V$} coefficient in Eq. (26). [^8] For a pure $\beta$-unstable species, that is no EC and thus $N_{0}$ only, we recover the DM formula given, for instance, in Maurin et al. (2022).
B.2 Formula for a single level ($n=1$)
Most (if not all) papers in the literature consider a single attachment only, i.e., zero or one electron attached ($n=1$ and $N_{2\penalty 10000\ {\rm lev}}=N_{0}+N_{1}$). In that case, the solution is
$\begin{cases}\displaystyle N_{0}=\frac{Q}{\alpha_{0}-\Gamma^{\rm att}\Gamma^{\rm str}\alpha_{1}^{-1}},\\ \displaystyle N_{1}=\frac{\Gamma^{\rm att}}{\alpha_{1}}N_{0}=\frac{\Gamma^{\rm att}}{\alpha_{1}}\times\frac{Q}{\alpha_{0}-\Gamma^{\rm att}\Gamma^{\rm str}\alpha_{1}^{-1}}.\end{cases}$ (30)
Substituting $\alpha_{i}$, $\cal P$}, and $\cal V$} from Eqs. (25) and (26) leads to
$\begin{cases}\displaystyle N_{0}^{\rm LBM}\!=\!\frac{Q}{(\tau_{\rm esc}\!+\!\Gamma^{\rm inel}\!+\!\Gamma^{\rm att})-\Gamma^{\rm att}\Gamma^{\rm str}\left(\tau_{\rm esc}^{-1}\!+\!\Gamma^{\rm inel}\!+\!\Gamma^{\rm str}\!+\!\Gamma^{\rm EC}/2\right)^{-1}},\\ \displaystyle N_{1}^{\rm LBM}=\frac{\Gamma^{\rm att}N_{0}^{\rm LB}}{\tau_{\rm esc}^{-1}+\Gamma^{\rm inel}+\Gamma^{\rm str}+\Gamma^{\rm EC}/2}.\end{cases}$ (31)
and
$\begin{cases}\displaystyle N_{0}^{\rm 1D-DM}\!=\!\frac{Q}{\left(\displaystyle\frac{D}{hL}\!+\!\Gamma^{\rm inel}\!+\!\Gamma^{\rm att}\right)-\frac{\displaystyle\Gamma^{\rm att}\Gamma^{\rm str}}{\left(\sqrt{\displaystyle\frac{D\,\Gamma^{\rm EC}}{2h^{2}}}\displaystyle\coth\!\left(\!\!\sqrt{\frac{L^{2}\Gamma^{\rm EC}}{2D}}\right)\!+\!\Gamma^{\rm inel}\!+\!\Gamma^{\rm str}\right)}},\\[42.67912pt] \displaystyle N_{1}^{\rm 1D-DM}=\frac{\Gamma^{\rm att}N_{0}^{\rm DM}}{\sqrt{\displaystyle\frac{D\,\Gamma^{\rm EC}}{2h^{2}}}\displaystyle\coth\left(\!\sqrt{\frac{L^{2}\Gamma^{\rm EC}}{2D}}\right)+\Gamma^{\rm inel}+\Gamma^{\rm str}}.\end{cases}$ (32)
Note that the latter solution slightly differs from the one given in Borchiellini et al. (2023), where the halving of the decay rate, $\tilde{\Gamma}^{\rm EC}=\Gamma^{\rm EC}/2$ in Eq. (22) for a single electron attached, was not accounted for.
Table 6: List of EC clocks isotopes (i.e., nuclei whose only relevant decay channels is EC) retrieved from NUBASE 2020 (Kondev et al., 2021). The columns show the EC clock (formatted as AXZAXZ with $X$ the element name), decay time, stable isotope at the end of its decay chain, and if applicable, the multistep decay chain. A few species with only upper limits on their unmeasured $\beta^{+}$ channels are also considered as pure EC-unstable: $^{56}{\rm Ni}$ with ${\cal B}r(\beta^{+})<$5.8\text{\times}{10}^{-05}$\$ ($t_{1/2}>$2.9\text{\times}{10}^{04}$\$ yr), $^{143}{\rm Pm}$ with ${\cal B}r(\beta^{+})<$5.7\text{\times}{10}^{-06}$\$ ($t_{1/2}>$1.3\text{\times}{10}^{07}$\$ yr), and $^{144}{\rm Pm}$ with ${\cal B}r(\beta^{+})<$8.0\text{\times}{10}^{-05}$\$ ($t_{1/2}>$1.2\text{\times}{10}^{06}$\$ yr).
EC clock $t_{1/2}$ [unit] End chain Multi-step decay path
4Be7 53.22 d 3Li7
18Ar37 35.011 d 17Cl37
20Ca41 99.4 kyr 19K41
22Ti44 59.1 yr 20Ca44 $^{44}{\rm Ti}$ $\rm EC$}, 59.1 yr ] $\to$ $^{44}{\rm Sc}$ $\beta^{+}$, 4.0421 h ] $\to$ $^{44}{\rm Ca}$
23V49 330 d 22Ti49
24Cr51 27.7015 d 23V51
25Mn53 3.7 Myr 24Cr53
26Fe55 2.7562 yr 25Mn55
28Ni56 6.075 d 26Fe56 $^{56}{\rm Ni}$ $\rm EC$}, 6.075 d ] $\to$ $^{56}{\rm Co}$ $\beta^{+}$, 77.236 d ] $\to$ $^{56}{\rm Fe}$
27Co57 271.811 d 26Fe57
28Ni59 81.82 kyr 27Co59
31Ga67 3.2617 d 30Zn67
32Ge68 271.05 d 30Zn68 $^{68}{\rm Ge}$ $\rm EC$}, 271.05 d ] $\to$ $^{68}{\rm Ga}$ $\beta^{+}$, 67.842 mn ] $\to$ $^{68}{\rm Zn}$
32Ge71 11.43 d 31Ga71
34Se72 8.40 d 32Ge72 $^{72}{\rm Se}$ $\rm EC$}, 8.40 d ] $\to$ $^{72}{\rm As}$ $\beta^{+}$, 26.0 h ] $\to$ $^{72}{\rm Ge}$
33As73 80.30 d 32Ge73
34Se75 119.78 d 33As75
36Kr81 229 kyr 35Br81
38Sr82 25.35 d 36Kr82 $^{82}{\rm Sr}$ $\rm EC$}, 25.35 d ] $\to$ $^{82}{\rm Rb}$ $\beta^{+}$, 1.2575 mn ] $\to$ $^{82}{\rm Kr}$
37Rb83 86.2 d 36Kr83
38Sr85 64.846 d 37Rb85
40Zr88 83.4 d 38Sr88 $^{88}{\rm Zr}$ $\rm EC$}, 83.4 d ] $\to$ $^{88}{\rm Y}$ $\beta^{+}$, 106.629 d ] $\to$ $^{88}{\rm Sr}$
42Mo93 4.0 kyr 41Nb93
43Tc97 4.21 Myr 42Mo97
46Pd100 3.63 d 44Ru100 $^{100}{\rm Pd}$ $\rm EC$}, 3.63 d ] $\to$ $^{100}{\rm Rh}$ ${\rm EC},\beta^{+}$, 1312.30,17.69 mn,d ] $\to$ $^{100}{\rm Ru}$
45Rh101 4.07 yr 44Ru101
46Pd103 16.991 d 45Rh103
48Cd109 461.3 d 47Ag109
50Sn110 4.154 h 48Cd110 $^{110}{\rm Sn}$ $\rm EC$}, 4.154 h ] $\to$ $^{110}{\rm In}$ $\beta^{+}$, 4.92 h ] $\to$ $^{110}{\rm Cd}$
49In111 2.8048 d 48Cd111
52Te118 6.00 d 50Sn118 $^{118}{\rm Te}$ $\rm EC$}, 6.00 d ] $\to$ $^{118}{\rm Sb}$ $\beta^{+}$, 3.6 mn ] $\to$ $^{118}{\rm Sn}$
51Sb119 38.19 h 50Sn119
54Xe122 20.1 h 52Te122 $^{122}{\rm Xe}$ $\rm EC$}, 20.1 h ] $\to$ $^{122}{\rm I}$ ${\rm EC},\beta^{+}$, 16.50,4.65 mn ] $\to$ $^{122}{\rm Te}$
53I125 59.392 d 52Te125
54Xe127 36.342 d 53I127
56Ba128 2.43 d 54Xe128 $^{128}{\rm Ba}$ $\rm EC$}, 2.43 d ] $\to$ $^{128}{\rm Cs}$ $\beta^{+}$, 3.640 mn ] $\to$ $^{128}{\rm Xe}$
55Cs131 9.689 d 54Xe131
56Ba133 10.5379 yr 55Cs133
58Ce134 3.16 d 56Ba134 $^{134}{\rm Ce}$ $\rm EC$}, 3.16 d ] $\to$ $^{134}{\rm La}$ $\beta^{+}$, 6.45 mn ] $\to$ $^{134}{\rm Ba}$
57La137 60 kyr 56Ba137
58Ce139 137.642 d 57La139
60Nd140 3.37 d 58Ce140 $^{140}{\rm Nd}$ $\rm EC$}, 3.37 d ] $\to$ $^{140}{\rm Pr}$ ${\rm EC},\beta^{+}$, 6.61,6.96 mn ] $\to$ $^{140}{\rm Ce}$
61Pm143 265 d 60Nd143
61Pm144 363 d 60Nd144
61Pm145 17.70 yr 60Nd145
62Sm145 340 d 60Nd145 $^{145}{\rm Sm}$ $\rm EC$}, 340 d ] $\to$ $^{145}{\rm Pm}$ $\rm EC$}, 17.70 yr ] $\to$ $^{145}{\rm Nd}$
64Gd146 48.27 d 60Nd142 $^{146}{\rm Gd}$ $\rm EC$}, 48.27 d ] $\to$ $^{146}{\rm Eu}$ $\beta^{+}$, 4.61 d ] $\to$ $^{146}{\rm Sm}$ $\alpha$, 68 Myr ] $\to$ $^{142}{\rm Nd}$
63Eu149 93.1 d 62Sm149
64Gd153 240.6 d 63Eu153
65Tb155 5.32 d 64Gd155
65Tb157 71 yr 64Gd157
68Er158 2.29 h 66Dy158 $^{158}{\rm Er}$ $\rm EC$}, 2.29 h ] $\to$ $^{158}{\rm Ho}$ $\beta^{+}$, 11.3 mn ] $\to$ $^{158}{\rm Dy}$
66Dy159 144.4 d 65Tb159
68Er160 28.58 h 66Dy160 $^{160}{\rm Er}$ $\rm EC$}, 28.58 h ] $\to$ $^{160}{\rm Ho}$ $\beta^{+}$, 25.6 mn ] $\to$ $^{160}{\rm Dy}$
67Ho161 2.48 h 66Dy161
67Ho163 4.570 kyr 66Dy163
70Yb164 75.8 mn 68Er164 $^{164}{\rm Yb}$ $\rm EC$}, 75.8 mn ] $\to$ $^{164}{\rm Tm}$ ${\rm EC},\beta^{+}$, 3.28,5.13 mn ] $\to$ $^{164}{\rm Er}$
68Er165 10.36 h 67Ho165
70Yb166 56.7 h 68Er166 $^{166}{\rm Yb}$ $\rm EC$}, 56.7 h ] $\to$ $^{166}{\rm Tm}$ $\beta^{+}$, 7.70 h ] $\to$ $^{166}{\rm Er}$
69Tm167 9.25 d 68Er167
70Yb169 32.014 d 69Tm169
72Hf170 16.01 h 70Yb170 $^{170}{\rm Hf}$ $\rm EC$}, 16.01 h ] $\to$ $^{170}{\rm Lu}$ $\beta^{+}$, 2.012 d ] $\to$ $^{170}{\rm Yb}$
72Hf172 1.87 yr 70Yb172 $^{172}{\rm Hf}$ $\rm EC$}, 1.87 yr ] $\to$ $^{172}{\rm Lu}$ $\beta^{+}$, 6.70 d ] $\to$ $^{172}{\rm Yb}$
71Lu173 1.37 yr 70Yb173
72Hf175 70.65 d 71Lu175
74W176 2.5 h 72Hf176 $^{176}{\rm W}$ $\rm EC$}, 2.5 h ] $\to$ $^{176}{\rm Ta}$ $\beta^{+}$, 8.09 h ] $\to$ $^{176}{\rm Hf}$
74W178 21.6 d 72Hf178 $^{178}{\rm W}$ $\rm EC$}, 21.6 d ] $\to$ $^{178}{\rm Ta}$ $\beta^{+}$, 2.36 h ] $\to$ $^{178}{\rm Hf}$
73Ta179 1.82 yr 72Hf179
74W181 120.956 d 73Ta181
76Os182 21.84 h 74W182 $^{182}{\rm Os}$ $\rm EC$}, 21.84 h ] $\to$ $^{182}{\rm Re}$ $\beta^{+}$, 64.2 h ] $\to$ $^{182}{\rm W}$
75Re183 70.0 d 74W183
76Os185 92.95 d 75Re185
77Ir189 13.2 d 76Os189
78Pt191 2.83 d 77Ir191
80Hg192 4.85 h 78Pt192 $^{192}{\rm Hg}$ $\rm EC$}, 4.85 h ] $\to$ $^{192}{\rm Au}$ $\beta^{+}$, 4.94 h ] $\to$ $^{192}{\rm Pt}$
78Pt193 50 yr 77Ir193
80Hg194 447 yr 78Pt194 $^{194}{\rm Hg}$ $\rm EC$}, 447 yr ] $\to$ $^{194}{\rm Au}$ $\beta^{+}$, 38.02 h ] $\to$ $^{194}{\rm Pt}$
79Au195 186.01 d 78Pt195
80Hg197 64.93 h 79Au197
82Pb200 21.5 h 80Hg200 $^{200}{\rm Pb}$ $\rm EC$}, 21.5 h ] $\to$ $^{200}{\rm Tl}$ $\beta^{+}$, 26.1 h ] $\to$ $^{200}{\rm Hg}$
81Tl201 3.0421 d 80Hg201
81Tl202 12.31 d 80Hg202
82Pb202 52.5 kyr 80Hg202 $^{202}{\rm Pb}$ $\rm EC$}, 52.5 kyr ] $\to$ $^{202}{\rm Tl}$ $\rm EC$}, 12.31 d ] $\to$ $^{202}{\rm Hg}$
82Pb203 51.924 h 81Tl203
82Pb205 17.0 Myr 81Tl205